OpenStudy (anonymous):

PLEASE HELP MEDAL AND FAN FOR FIRST/BEST ANSWER!!! Melissa plays on her school's bolleyball team. She needs to fill up 4 volleyballs with air. Each volleyball has a radius of 12 cm. How much air, IN TOTAL, will it take to fill up all the volleyballs?

OpenStudy (anonymous):

each one has volume \[\frac{4}{3}\pi r^3\] and in our case \(r=12\) so each one has volume \[\frac{4}{3}\pi 12^3\]

OpenStudy (anonymous):

since there are four of them, compute \[4\times \frac{4}{3}\pi \times 12^3\]

OpenStudy (anonymous):

im confused. how do you even solve an equation like that? cant you just find the volume of one volleyball and then multiply that by 4?

OpenStudy (anonymous):

@satellite73

OpenStudy (anonymous):

which would make the answer 45,976.36 right???

OpenStudy (anonymous):

@mathman123 @satellite73